# Perception-as-Control: Fine-grained Controllable Image Animation with 3D-aware Motion Representation

### v1

@@ 3.3 Network Architecture
We propose a motion-controllable image animation framework called Perception-as-Control based on 3D-aware motion representation, which takes a reference image and a sequence of spatially-aligned motion control signals as input and generates a video as output.
As shown in the right part of Fig. 3, our framework is based on a denoising U-Net architecture, with extra motion modules [8] for modeling temporal information. Under the image animation setting, it is essential to preserve the appearance information contained in the reference image (image control signals) and follow the motion information contained in motion control signals. Previous methods inject spatial-aligned control signals by directly adding them to the input noise [19, 22] or integrating them into denoising U-Net through auxiliary architectures [44, 12, 23], and inject high-level control signals via attention mechanism [41, 21]. Given that the motion control signals are already spatially aligned, in our framework, we inject motion control signals by adding them to the noise and inject the reference image via ReferenceNet [12]. Specifically, we use two lightweight encoders to encode camera control signals and object control signals separately, avoiding RGB-level interference. Then, a fusion module merges the encoded camera control signals and object control signals. We implement the fusion module empirically as a convolution block.
During training, pairs of training data are prepared through our data curation pipeline, and a three-stage training strategy is applied for balance camera and object control signals. During inference, user intentions conveyed in various forms are transformed into 3D-aware motion representation, and the proposed framework is used to generate videos based on the representation.
@@ 3.4 Training Strategy
Although the proposed 3D-aware motion representation transforms camera and object motion into a unified form, the training strategy is still important since motions in training data are entangled. To this end, we introduce our three-stage training strategy, which aims to balance camera and object control signals for achieving fine-grained collaborative motion control.
Stage 1: Camera Motion Control Training.
In the first stage, we use video clips from RealEstate10K[48] with camera motion only to train our camera encoder. We temporarily exclude object encoder and fusion module from the framework. Since the rendered spheres can be anywhere in the pixel plane, motion modules in denoising U-Net are trained together. This allows the model to align each rendered sphere with its represented object parts. Thus, we train camera encoder, motion modules in denoising U-Net, and ReferenceNet in this stage.
Stage 2: Collaborative Motion Control Training.
In the second stage, we add video clips from WebVid10M[1] that contain both camera and object motions for training. We add object encoder and fusion module back for collaborative motion control. To speed up the alignment between unit spheres and object parts, we utilize dense unit spheres as object control signals in this stage. We fix camera encoder and train object encoder, fusion module, motion modules, and ReferenceNet in this stage. Moreover, to preserve the effectiveness of both control signals simultaneously, especially in conflict cases, we randomly drop one type of control signals at a fixed dropout rate for video clips from RealEstate10K during training.
Stage 3: Dense-to-Sparse Fine-tuning.
The goal of the third stage is to achieve fine-grained object motion control via sparse unit spheres, which requires the model to determine the control range of each rendered point adaptively.
Considering that object motion tends to occur on salient objects in a video clip, we employ an off-the-shelf segmentation algorithm [17] to obtain salient objects in the reference image and select unit spheres inside the segmentation mask as Set1. Then we calculate the trajectory length of each unit sphere and select those with a trajectory length higher than the $80^{\rm th}$ percentile as Set2. By combining Set1 and Set2, we obtain the dense sphere set and randomly sample $N\in\{1,2,\dots,16\}$ spheres from it as the final sparse sphere set per iteration. In this stage, we fix all components except fusion module and motion modules during training.
In all stages, we optimize our framework with a loss function similar to Eq. 2:
@@
where $\epsilon\sim N(0,1)$ is drawn from standard Gaussian noise, $t\in{1,\dots,T}$ denotes the diffusion timestep, and $c_{\rm img}$, $c_{\rm cam}$, $c_{\rm obj}$ are the reference image, camera motion control signals and object motion control signals, respectively.
Figure 4: Case study of camera-only motion control.
@@ 4.1 Experimental Settings
We use video datasets RealEstate10K [48] and WebVid10M [1] for training. RealEstate10K contains videos with only camera motion, and each video is annotated with ground-truth camera intrinsic and extrinsic parameters. WebVid10M is a large-scale dataset containing in-the-wild videos with both camera and object motions, while the quality of these videos is not high enough. After our data curation pipeline, our final training data contains around 6K video clips from RealEstate10K and 35K video clips from WebVid10M.
Implementation Details.
@@
We adopt Stable Diffusion 1.5 (SD1.5) [26] as our base model, and initialize the weights of both denoising U-Net and ReferenceNet using the pretrained weights of SD1.5. We add motion modules the same way as previous works [8, 2] with random initialization. During training, the weights of VAE [14] encoder and decoder, as well as the CLIP [25] image encoder are frozen. During training, we randomly sample 16 frames for each video clip with a resolution of $768\times 512$. We use AdamW [18] as an optimizer. We conduct the training on 8 NVIDIA A100 GPUs. We train around 20k iterations for stage 1, 20k iterations for stage 2, and 50k iterations for stage 3, with a batch size of 1 for convergence.
Figure 5: Case study of object-only motion control.
Figure 6: Case study of collaborative motion control.
4.2 Motion-controllable Image Animation
4.2.1 Camera-only Motion Control
To demonstrate the superiority of Perception-as-Control in camera-only motion control, we select both basic and arbitrary camera motions and visualize them in 3D. As shown in Fig. 4, our framework can generate videos that adhere to the expected camera motion. Owing to the world envelope that shows the direction and magnitude of camera motion and enhances the overall perception, the spatial structures in the generated videos are stable and consistent. Moreover, by adding several unit spheres to essential parts, the generated videos are aware of depth information and can accurately reflect perspective effects under different focal lengths.
Figure 7: Various motion-related applications. Red boxes denote reference images or source videos.
4.2.2 Object-only Motion Control
For object-only motion control, we provide cases in Fig. 5 to demonstrate the superiority of our method in achieving flexible and precise multi-instance and fine-grained motion control. Since there is no camera motion, we project unit sphere centers onto the pixel plane and visualize how they change over time, using colors to represent the direction of movement. As shown in the left part of Fig. 5, we control an increasing number of hot air balloons in the reference image, and the generated videos precisely reflect the movements of each one. In the right part of Fig. 5, when additional controls are applied to the bamboo, the proposed method can adaptively determine the control range for each unit sphere.
4.2.3 Collaborative Motion Control
Our framework also supports collaborative control of both camera and object motions. As illustrated in Fig. 6, we control both camera and object motions and visualize the corresponding 3D-aware motion representation. Our method can handle both large object motions and delicate object motions. The generated videos adhere to the specified camera motion and accurately present adapted object motions.
@@ 4.3 Applications
The proposed image animation framework can be used for various motion-related applications by carefully interpreting user intentions, as illustrated in Fig. 1.
Motion Generation refers to image animation based on a reference image and 2D trajectories drawn by users. Motion Clone refers to mimicking both camera and object motions in a source video. Users provide a source video to construct 3D-aware motion representation and edit the first frame as the reference image. Motion Transfer refers to transferring local motions from the source video to the reference image by adaptively relocating and scaling them according to semantic correspondence. Motion Editing has greater control freedom than the above settings. Users can provide segmentation masks and edit motions inside or outside the segmentation masks.
More examples are shown in Fig. 7. For motion generation, we construct 3D-aware motion representation based on the reference image via monouclar depth estimation and manipulate it according to 2D trajectories. The results show that our framework effectively handles various motions including natural swaying, shape changes, intersections, and movements that extend out of the pixel plane. For motion clone, we edit the first frame of each source video as the reference image and construct 3D-aware motion representation from the source video as control signals. The results exactly reproduce the camera and object motions in the source video. For motion transfer, we extract 3D trajectories of unit spheres corresponding to semantic points and relocate them to match the reference image. The results show that local motions from the source video are successfully transferred to objects of different scales at different locations in the reference images. For motion editing, we modify 3D-aware motion representation from the source video based on segmentation masks, e.g., modify several unit sphere trajectories or viewpoints. The results demonstrate the flexibility of our framework.
@@ 5 Conclusion
Unlike previous methods that only control camera motion or object motion, we propose a fine-grained controllable image animation framework for collaborative motion control. The framework takes motion control signals derived from 3D-aware motion representation as input, simplifying the original 3D scene to transform camera and object motions into unified and consistent visual changes. Due to the flexibility of the introduced 3D-aware motion representation, our framework can support various motion-related applications by translating user intentions as changes in the representation.

### v2

@@ 3.3 Network Architecture
We propose an image animation framework to incorporate our 3D-aware motion representation, which takes a reference image and a sequence of spatially aligned motion control signals as input and generates a video as output.
As shown in Fig. 3, our framework is based on a denoising U-Net architecture, with extra motion modules [7] for modeling temporal information. Under the image animation setting, it is essential to preserve the appearance information contained in the reference image (image control signals) and follow the motion information contained in motion control signals. Previous methods inject control signals by directly adding them to the input noise [21, 24] or integrating them into denoising U-Net through auxiliary architectures [48, 45, 23, 13, 37, 9, 25]. Given that the motion control signals are already spatially aligned, in our framework, we inject motion control signals by adding them to the noise and inject the reference image via ReferenceNet [13] for better appearance preservation. Specifically, we use two lightweight encoders to encode camera and object control signals separately, avoiding RGB-level interference, and then use a fusion module to merge them as the final motion control signals. We implement the fusion module empirically as a convolution block.
During training, training data are prepared through our data curation pipeline, and a three-stage training strategy is applied for balancing camera and object motion control. During inference, user intentions conveyed in various forms are transformed into motion control signals via 3D-aware motion representation, as the input of our framework.
@@ 3.4 Training Strategy
Although both camera and object motions can be transformed into unified and consistent visual changes through 3D-aware motion representation, video clips with entangled motions increase the difficulty of training.
Stage 1: Camera Motion Control Training. In the first stage, we use video clips from RealEstate-10K[52] with camera motion only to train our camera encoder. We temporarily exclude object encoder and fusion module from the framework. Since pure viewpoint changes bring about coherent and consistent overall scene movements, we train motion modules together to capture the correspondence between changes of the world envelope and scene contents. Thus, we optimize camera encoder, motion modules, and ReferenceNet in this stage.
Stage 2: Collaborative Motion Control Training. In the second stage, we add video clips from WebVid-10M[1] that contain both camera and object motions for training. We add object encoder and fusion module back for collaborative motion control. Motion modules are optimized to align each rendered unit sphere with its corresponding object parts, since the initial location of each unit sphere is not restricted. To speed up the alignment between unit spheres and object parts, we use dense unit spheres in this stage. We fix camera encoder and optimize object encoder, fusion module, motion modules, and ReferenceNet. Moreover, to preserve the effectiveness of both control signals, especially in conflict cases, we randomly drop one type of control signal at a fixed rate for RealEstate10K video clips.
Stage 3: Dense-to-Sparse Fine-tuning. This stage aims to achieve fine-grained motion control with our world envelope and sparse unit spheres, which requires the model to determine each sphere’s control range adaptively. Considering that object motion tends to occur on salient objects, we employ an off-the-shelf segmentation algorithm [19] to obtain salient objects in the reference image and select unit spheres inside their mask as Set1. Then we calculate the trajectory length of each unit sphere and select those with a trajectory length higher than the $80^{\rm th}$ percentile as Set2. By combining Set1 and Set2, we obtain the dense sphere set and randomly sample $N\in\{1,2,\dots,16\}$ spheres from it per iteration. In this stage, we fix all components except fusion module and motion modules during training.
For all stages, the loss function is in Eq. 3:
@@
where $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is drawn from standard Gaussian noise, $t\in\{1,\dots,T\}$ denotes the diffusion timestep, and $c_{\rm img}$, $c_{\rm cam}$, $c_{\rm obj}$ are the reference image, camera motion control signals and object motion control signals, respectively.
Figure 4: Results of fine-grained collaborative motion control.
@@ 4 Experiments
[…]
CameraCtrl 14.538 178.497 0.648 0.203
Ours14 11.795 73.449 0.641 0.199
MotionCtrl 17.081 122.570 0.957 0.278
CameraCtrl* 14.942 136.602 0.837 0.257
MotionCtrl++ 16.327 104.020 0.906 0.265
CameraCtrl*++ 12.923 113.165 0.761 0.291
Ours 9.675 52.421 0.757 0.235
Table 1: Comparison on RealEstate test set for camera motion control. CameraCtrl* denotes the 16-frame version we reproduced.
[…]
Motion-I2V 60.643 367.524 34.52 0.845
MOFA-Video 26.895 175.701 30.26 0.944
MotionCtrl++ 32.528 208.387 28.57 0.959
CameraCtrl*++ 28.393 164.654 25.72 0.963
Ours 16.465 161.076 23.32 0.968
Table 2: Comparison on WebVid test set for object motion control.
Figure 5: Qualitative comparison with state-of-the-art methods regarding camera motion control. Reference images are generated by reconstructing point clouds based on the reference image and rendering them according to camera parameters.
Figure 6: Qualitative comparison with state-of-the-art methods regarding object motion control and simple collaborative motion control.
@@ 4.1 Experimental Settings
Datasets. We use video datasets RealEstate-10K [52] and WebVid-10M [1] for training. RealEstate-10K contains videos with pure camera motion, and each video is annotated with ground-truth camera parameters. WebVid-10M is a large-scale dataset containing in-the-wild videos with both camera and object motions. After our data curation pipeline, the training data contains around 61K video clips from RealEstate and 352K from WebVid.
Metrics. We adopt Frechet Video Distance (FVD) [34], Frechet Inception Distance (FID) [10] to measure visual quality. ObjMC [37] and Subject Consistency (SubCon) in VBench [14] are used to measure object motion controllability and average rotation error (RotErr) and translation error (TransErr) [9] are used to measure camera motion controllability. We select 1K clips from RealEstate test set and 1K from WebVid test set for quantitative evaluation.
@@
Implementation Details. We adopt Stable Diffusion 1.5 (SD1.5) [28] as our base model, and initialize both denoising U-Net and ReferenceNet [13] using its pretrained weights. We add motion modules as in previous works [7, 2] with random initialization. During training, CLIP [27] image encoder and VAE [16] encoder and decoder remain frozen. The model generates videos with 16 frames at a resolution of $768\times 512$. We use AdamW [20] as optimizer. We conduct the training on one NVIDIA A100 GPU. We train around 20k iterations for both stage 1 and stage 2 with learning rate of 1e-5, and 50k for stage 3 with learning rate of 1e-6, with a batch size of 1 for convergence.
Figure 7: Qualitative evidence for ablation study.
[…]
2D point offset 26.823 254.907 53.28 0.967
2D point position 16.520 177.710 30.61 0.968
$Sph$ (3D) 16.458 176.989 29.47 0.968
$Env$ + $Sph$ (Ours) 16.465 161.076 23.32 0.969
Table 3: Ablation study on 3D-aware motion representation. $Env$ denotes our world envelope and $Sph$ denotes 3D unit spheres.
[…]
One-stage 59.040 318.941 36.72 0.959
Two-stage 26.823 254.907 31.29 0.966
Three-stage (Ours) 16.465 161.076 23.32 0.969
Table 4: Ablation study on the proposed training strategy.
Figure 8: Results on four kinds of motion-related applications. More details and results please refer to Supplementary Materials.
4.2 Comparison with State-of-the-Art Methods
We compare our approach with state-of-the-art methods in Tab. 1 and Tab. 2. We use the SVD version of MotionCtrl [37] and CameraCtrl [9] for image-based generation. MotionCtrl directly takes camera extrinsic parameters as control signals, and CameraCtrl uses Plücker embeddings. Motion-I2V [29] and MOFA-Video [25] employ dense optical flow as the final control signals. Since the proposed 3D-aware motion representation is orthogonal to specific base models, we also incorporate our representation into CameraCtrl and MotionCtrl and train them following our training strategy to empower them with fine-grained collaborative motion control capability, denoted as CameraCtrl++ and MotionCtrl++, respectively (see Sec. B in Supp.). By incorporating our 3D-aware motion representation in CameraCtrl and MotionCtrl, their ++ versions perform better on camera motion control and are capable of collaborative motion control. Moreover, ours outperforms their ++ versions, which shows the superiority of our framework.
Our approach achieves the best performance in terms of both camera and object motion control, as further evidenced in Fig. 5 and Fig. 6. In Fig. 5, we notice that CameraCtrl degrades on out-of-distribution camera trajectories such as large orbitary ones and tends to lose details in later frames, and MotionCtrl fails to adhere to the designated camera trajectories strictly and causes image blurring under large camera motions. In contrast, our approach generalizes well and achieves precise camera motion control. From Fig. 6 we can observe that optical flow-based methods face conflict control issues in collaborative motion control and struggle to maintain object appearance under large motions.
4.3 Ablation Study
We conduct ablation studies to demonstrate the effectiveness of 3D-aware motion representation and validate the necessity of our three-stage training strategy. In Tab. 3, 2D point offset refers to motion representations in previous works [37, 18], and 2D point position is similar to visualizing the projection of unit sphere centroids with a fixed radius. We have these observations: 1) 2D point offset underperforms others, and we argue that this form is neither as complete as dense optical flow nor visually aligned as point positions, making it hard to associate points with object parts in the reference image. 2) The lack of depth information in 2D point position makes the spatial relations less clear and harms performance (Fig. 7a). 3) Only using 3D unit spheres may cause the ambiguity problem (Fig. 7b).
In Tab. 4. One-stage refers to directly optimizing all the modules mentioned using mixed data and Two-stage refers to merging the last two stages into one. The results show that our carefully designed training strategy is essential in reducing training difficulty and achieving harmonious collaborative motion control (See Sec. C in Supp.).
4.4 Fine-grained Collaborative Motion Control
Fig. 4 shows the superiority of Perception-as-Control on fine-grained collaborative motion control. For clarity, we project unit sphere centroids onto the reference image to visualize how they change over time (colors represent movement direction). In the first row, we set object motions for several hot air balloons, and the result precisely reflects each one. In the second row, when setting several unit spheres to the bamboo simultaneously, our approach adaptively determines the control range for each one. In the last two rows, we set both camera and object motions, the results adhere to the camera motion and present the adapted object motions (see Sec. A in Supp.).
@@ 4.5 Applications
As illustrated in Fig. 1, our approach has the potential to support various motion-related applications. More results are given in Fig. 8. For motion generation, we construct 3D-aware motion representation based on the reference image and manipulate it according to user-provided 2D/3D trajectories. Our approach handles motions including natural swaying, shape changes, intersections, and large displacements. For motion clone, we construct 3D-aware motion representation from the source video as motion control signals, and edit the first frame as the reference image. The results exactly mimic the entire motions (camera and object) in the source video. For motion transfer, we extract 3D trajectories of unit spheres corresponding to semantic points and relocate them to match those in the reference image. The results show that local motions from the source video are successfully transferred to objects that differ in scale and location. For motion editing, we partially manipulate the 3D-aware motion representation based on user-specified regions, e.g., reset unit sphere trajectories or viewpoint changes. The results demonstrate the flexibility of our framework (see Sec. D in Supp.).
@@ 5 Conclusion
This paper aims to achieve precise and flexible fine-grained collaborative motion control, unlike previous methods that only support coarse-grained or inflexible collaborative motion control. We observe the defects of previous motion representations and tackle the problems by proposing 3D-aware motion representation. This representation simplifies the corresponding 3D scene by transforming camera and object motions into unified and consistent visual changes. The proposed image animation framework takes motion control signals derived from 3D-aware motion representation and can support various motion-related applications. Our extensive experiments show that our 3D-aware motion representation is orthogonal to model architectures and improves motion controllability consistently on multiple benchmarks (see Sec. E in Supp. for limitations).
Supplementary Material
Figure 9: Results of camera-only motion control using our Perception-as-Control.
Figure 10: Results with the same camera extrinsic sequence (Pedestal Up) under different focal lengths.
A. More Results of Perception-as-Control
Perception-as-Control can achieve precise camera motion control adhering to any designated camera trajectory, multi-instance object motion control with adaptive control granularity, and harmonious collaborative motion control.
Figure 11: Results of object-only motion control using our Perception-as-Control.
A.1. Camera-only Motion Control
To demonstrate the superiority of Perception-as-Control in camera-only motion control, we show the generated results under different camera trajectories in Fig. 9. Our approach places no restrictions on the camera trajectories and supports camera motion control under basic camera trajectories such as Pan Left and Dolly In, their combinations such as Pan Left + Zoom In, and any user-provided camera trajectories. Please refer to our webpage: Fine-grained Collaborative Motion Control - Camera-only Motion Control for the generated videos.
As shown in Fig. 9, our framework can generate videos that adhere to any designated camera trajectories. The generated videos maintain stable and consistent spatial structures owing to the world envelope in our 3D-aware motion representation, which demonstrates the direction and magnitude of camera movements and enhances overall perception. Moreover, by adding several unit spheres to essential object parts, the generated videos are aware of depth information and can accurately reflect perspective effects under different focal lengths, as shown in Fig. 10.
A.2. Object-only Motion Control
For object-only motion control, we show the generated results in Fig. 11 to demonstrate the superiority of our approach in achieving flexible and precise multi-instance object motion control with adaptive control granularity. Please refer to our webpage: Fine-grained Collaborative Motion Control - Object-only Motion Control.
For visualization clarity, we project the unit sphere centroids onto the reference image to show how they change over time, using colors to represent the direction of movement. As shown in the left part of Fig. 11, we control an increasing number of hot air balloons in the reference image, and the generated videos precisely reflect the movements of each one. In the right part of Fig. 11, when additional controls are applied to the bamboo, the proposed method can adaptively determine the control range for each unit sphere and achieve harmonious collaborative control results.
Figure 12: Results of collaborative motion control.
A.3. Collaborative Motion Control
Our framework also supports collaborative control of both camera and object motions. Please refer to our webpage: Fine-grained Collaborative Motion Control - Collaborative Motion Control for the generated videos.
As illustrated in Fig. 12, we control both camera and object motions and visualize the corresponding 3D-aware motion representation. Our method can handle both large object motions and delicate object motions. The generated videos adhere to the specified camera motion and accurately present the adapted object motions.
B. More Details of ++ Versions
To demonstrate that our 3D-aware motion representation is orthogonal to base model architectures and effective for different base model architectures, we incorporate 3D-aware motion representation into the SVD [2] version of CameraCtrl [9] and MotionCtrl [37] to empower them with fine-grained collaborative motion control capability. Without making major changes to their architectures, we successfully incorporate our motion representation into them, as shown in Fig. 13. The experimental results verify the effectiveness of our 3D-aware motion representation and the superiority of our framework.
Figure 13: The frameworks of MotionCtrl++ and CameraCtrl++.
C. More Details of Ablation Study
In Fig. 14, we illustrate the three training strategies included in our ablation study (Table 4). One-stage refers to directly optimizing all the modules mentioned using mixed data and Two-stage refers to merging the last two stages into one.
Figure 14: Illustration of different training strategies in ablation study.
D. More Details of Applications
Figure 15: Illustration of different motion-related applications.
By carefully interpreting user intentions into manipulation on 3D-aware motion representations, our framework can support various potential motion-related applications. In this paper, we describe four kinds of motion-related applications, including motion generation, motion clone, motion transfer, and motion editing. The intuitive illustration is shown in Fig. 15. Please refer to our webpage: Potential Applications for the generated videos.
Motion Generation refers to the task setting of interactive image animation that allows users to draw 2D or 3D point trajectories on the reference image and set an arbitrary camera trajectory with a carefully designed user interface. In this application, we employ an off-the-shelf metric depth estimation algorithm on the reference image, construct our 3D-aware motion representation based on the estimated results. We manipulate unit spheres corresponding to the starting point of each trajectory, and perceive the representation based on user-designated camera trajectory. The perception results are used as motion control signals. Ideally, the object parts corresponding to the starting point of each trajectory should move along the trajectory in the generated video. The difficulty lies in correctly understanding the control range of each trajectory point and generating reasonable and accurate motion videos under the control of multiple trajectories.
Motion Clone refers to extracting all motions from a source video and cloning them for driving a reference image. Users begin by providing a source video, which is used to construct our 3D-aware motion representation. They then edit the first frame of the source video to create a reference image. After that, we use the constructed 3D-aware motion representation and the reference image as inputs of the proposed framework to obtain the motion clone results.
Motion Transfer, similar to motion clone, refers to transferring local object motions from the source video to the reference image. In this case, the object may differ from that in the source video in terms of scale and location. To meet the needs of this application, we first find the semantic correspondences between the first frame of the source video and the reference image. Then, we construct our 3D-aware motion representation from the source video and locate essential unit spheres based on semantic correspondence points on the first frame. after that, we adaptively relocate and rescale them to match those semantic correspondence points on the reference image. Finally, the manipulated 3D-aware motion representation can be used to animate the reference image.
Motion Editing has greater control freedom than the above settings. Based on their source video, users provide segmentation masks and instructions to edit motions inside or outside these masks, i.e., editing fine-grained local motions in user-specified regions. In this way, users could freeze or modify motions of several objects or the background, etc. To achieve such operations, we first construct 3D-aware motion representation from the source video, and select unit spheres inside the regions to be edited. Then, we manipulate those selected unit spheres according to user intentions and use the manipulated representation to re-animate the first frame of the source video.
Figure 16: Limitations of Perception-as-Control.
E. Limitations and Future Work
Our work introduces a novel 3D-aware motion representation for motion-controllable image animation task. Although our image animation framework that incorporates 3D-aware motion representation achieves promising and insightful results in fine-grained collaborative motion control, it still has some limitations.
First, we simplify object parts as unit spheres, which can hardly reflect object rotation. As shown in the first row of Fig. 16, the relative positions between the two billiards and the markers on them remain unchanged, i.e., those billiards only move in translation without rotation. In future work, exploring more powerful alternatives may help express more comprehensive forms of movement.
Second, our approach is constrained by base model performance and cannot handle human-related motions satisfactorily. As shown in the second row of Fig. 16, the boy is moving forward without reasonable stepping movements. It is promising to incorporate the proposed 3D-aware motion representation into more powerful base model architectures.
Also, those off-the-shelf algorithms for constructing our 3D-aware representations influence the performance of our approach. Future work could employ state-of-the-art deep visual SLAM works for more precise estimation.